# On Reducing the Execution Latency of Superconducting Quantum Processors via Quantum Program Scheduling

### v1

@@ 3.2.2. Qubit Partitioning
After rearranging the updated queue, we need to select a number of programs to be executed in parallel in this iteration. Selected programs must share no common physical qubits with each other, so we should partition the coupling graph into separate parts. Concretely, we pick out the first program in the sorted queue to conduct qubit partitioning. If the partitioning algorithm cannot find a valid partition, it will skip the program, and proceed to the next. Otherwise, we will go on to the qubit mapping step. We use the qubit partitioning algorithm introduced in (Niu and Todri-Sanial, 2023). This method choose physical qubits with higher degrees than the largest logical degree as starting points. If such qubits do not exist, it chooses physical qubits with the highest degree as starting points. Then, it adds a neighboring qubit with the highest fidelity degree to the partition iteratively until the number of selected physical qubits equals that of logical qubits. Finally, the partition with the best fidelity score is selected.
@@ 3.2.3. Qubit Mapping
(3) $EPST^{*}=\prod_{i=1}^{N_{1q}}r_{1q}^{(o_{i})}\cdot\prod_{i=1}^{N_{2q}}r_{2q}^{(c_{i})}\cdot\prod_{i=1}^{N_{ro}}r_{ro}^{(m_{i})},$
where $r_{1q}^{(*)}$, $r_{2q}^{(*)}$, and $r_{ro}^{(*)}$ denote the reliability of one-qubit gates, two-qubit gates, and measurements. $N_{1q}$, $N_{2q}$, and $N_{ro}$ is the number of the three operations. $o_{i}$, $c_{i}$, and $m_{i}$ map operations to their locations in the circuit. Our $EPST^{*}$ score is more accurate than $EPST$, especially given high variance of reliability. We integrate $EPST^{*}$ in our noise-aware initial mapping algorithm in Alg. 1
@@ 11             $Best\_initial\_mapping=Initial\_mapping$
3.3. Greedy Baselines for Comparison
As there are no existing methods addressing the scheduling problem in this paper, we also devise three greedy methods as baselines. We regard the number of qubits $n$, number of shots $s$, and submission time $t$ as three most important factors for the time and fidelity metric in QPSP. Hence, we sort programs by these three properties in ascending order in the queue rearranging step to construct three greedy baselines, i.e. Smallest Number of Qubits First (SNQF), Smallest Number of Shots First (SNSF), and Earliest Submission Time First (ESTF). The reason is as follows:
SNQF. When $n$ is small, the QPU can accommodate more programs, which means more programs are executed in unit time. This is similar to the Shortest Job First (SJF) strategy in process scheduling, which significantly raises the throughput at the beginning, thus improving the average turnaround time.
SNSF. Since the shot number is set as the maximum of $s$ among programs in one execution, sorting by $s$ can decrease the number of unnecessary shots, which in turn reduces the QPU time.
ESTF. This method keeps the arriving order of jobs, sacrificing the turnaround time and QPU time for fairness, which is embodied in the maximum and standard deviation of turnaround time.
Note that our method can be degraded to these greedy queue rearranging if needed. For example, by setting $\alpha=1$, $\beta=0$, and $\gamma=0$, we get the greedy queue rearranging of SNQF. For baselines, we do not restrain the maximum usage of physical qubits, but execute the mapped programs as long as the current program cannot be mapped on the QPU by tradition.
Table 1. Performance comparison between different methods on the noise model and Xiaohong ( QuantumCTek).
@@ 3.3. Greedy Baselines for Comparison
Noise Model Default 919.48 0 5148 2587 0 1483 0 108 1 71.84 0
Ours 442.60 -51.86 2181 800 -69.06 584 -60.61 112 2.41 69.37 -2.47
SNQF 503.13 -45.28 2164 820 -68.29 596 -59.81 119 2.40 69.27 -2.57
SNSF 386.13 -58.01 2159 952 -63.20 640 -56.84 111 2.43 69.31 -2.54
ESTF 499.41 -45.69 2132 1070 -58.62 609 -58.91 110 2.43 69.45 -2.39
QuMC 730.71 -20.53 3810 2011 -22.25 1040 -29.86 215 1.42 71.92 0.07
Xiaohong Default 467.58 0 4728 2378 0 1361 0 89 1 39.74 0
Ours 85.27 -81.76 678 162 -93.20 173 -87.31 86 7.18 34.57 -5.17
SNQF 108.90 -76.71 593 181 -92.39 184 -86.47 87 6.85 35.62 -4.12
SNSF 71.79 -84.65 659 235 -90.14 195 -85.70 93 7.42 33.72 -6.02
ESTF 105.79 -77.38 533 268 -88.73 143 -89.52 94 7.40 34.00 -5.74
QuMC 293.08 -37.32 2544 964 -59.47 735 -46.00 584 1.98 35.93 -3.81
@@
  • $\Delta$ QPU Time(%): percentage difference to the QPU time of Default. TAT[s]: turnaround time in seconds. RT[s]: runtime of scheduling algorithms in seconds. $\Delta$ PST[%]: difference to PST of default in percent. $\Delta$ avg(%): percentage difference to the average of Default. $\Delta$ std(%): percentage difference to the standard deviation of Default.
3.4. Complexity Analysis for All the Methods
Given the number of programs $K$, the number of gates $g$, the number of physical qubits $N$, the number of logical qubits $n$, the number of starting points $m$ ($m<N$), and the number of repeats $r$, we can calculate the time complexity of our method and greedy baselines. The complexity of qubit partitioning and routing is $O\big{(}mn^{2}+N\log(N)+g\big{)}$ and $O(gN^{2.5})$, respectively (Niu and Todri-Sanial, 2023).
Queue rearranging. The main overhead of queue rearranging lies in sorting the queue, which takes $O\big{(}K\log(K)\big{)}$ time. Hence, the complexity of queue arranging is $O\big{(}K\log(K)\big{)}$.
Since every program should undergo qubit partitioning, initial mapping, and routing, the total complexity is $O(KrgN^{2.5}+Krn+Kmn^{2})$. Therefore, the overall time complexity is $O\big{(}K\log(K)+KrgN^{2.5}+Krn+Kmn^{2}\big{)}$. In normal circumstances, the number of repeats $r$ is a small constant, so the complexity can be reduced to $O\big{(}K\log(K)+KgN^{2.5}+Kmn^{2}\big{)}$. The routing overhead $O(KgN^{2.5})$ is the dominant part, which is unavoidable. Queue rearranging only incurs trivial overhead compared with routing.
4. Experiments
4.1. Protocols
4.1.1. Dataset
We construct our dataset from RevLib (Soeken et al., 2012), a benchmark of reversible and quantum circuits, which is widely used in related works (Li et al., 2019; Liu and Dou, 2021; Niu and Todri-Sanial, 2023). Circuits with large width or depth are unsuitable for either the noise model or real quantum hardware, so we filter the data. First, we choose the circuits with width smaller than 17 to be suitable for the 16-qubit noise model. Second, we translate circuits to fit the basis gate set of the noise model and Xiaohong respectively. Third, we choose the translated circuits with depth smaller than 100 to guarantee relatively high fidelity. Finally, the number of candidate circuits on the noise model and Xiaohong is 77 and 20, respectively. The range of width for them is $[3,16]$ (6 and 6.65 on average). The range of depth is $[5,99]$ and $[21,98]$ (44.1 and 64.5 on average).
Then, we sample from candidate circuits to construct our dataset. We focus on the congestion scene, where there are some initial jobs and much more jobs to be submitted. Due to the limitation of quantum resources and time, the number of initial jobs is 44 on average, and the number of new submitted jobs is 400. The submission time $t_{i}$ of initial jobs is set as 0. For new submitted jobs, $t_{i+1}\in\{t_{i},t_{i}+1\}$.
According to our observation, at peak periods on IBM quantum cloud, there are approximately two jobs submitted per second on average. Hence, the ratio is in line with the congestion in reality. For the noise model, the number of shots $s_{i}$ in each job ($p_{i},s_{i},t_{i}$) is set as a random integer from 1K to 20K. For Xiaohong (QuantumCTek), we modify the range of $s_{i}$ as $[500,10K]$ to reduce running overhead. The length of dataset is 6.
4.1.2. Baselines
To verify the effectiveness of our method, we compare it with the current running mode of quantum computers, which we denote as Default. Specifically, each submitted quantum program is executed in serial according to their submission time. In addition, we also include all our three greedy methods introduced in Sec. 3.3, and adapt QuMC (Niu and Todri-Sanial, 2023) to fit QPSP as baselines.
4.1.3. Parameter Setting and Experiment Environment
@@
Figure 4. (a) Impact of $\eta$ on PST and average TAT on Xiaohong. (b) Impact of $\gamma$ on average TAT and standard deviation of TAT on the noise model. (c) Impact of $\alpha$ and $\beta$ on average TAT on the noise model ($\gamma=0.5$). (d) Impact of $\alpha$ and $\beta$ on QPU time on the noise model ($\gamma=0.5$). The blue bar in (c) and (d) corresponds with the hyperparameters we pick for the noise model.
4.2. Results on a Simulated Noise Model
@@
As shown in Tbl. 1, our method achieves the shortest average turnaround time (TAT) across all methods. It reduces TAT of Default by nearly 70 percent, which will significantly cut down the waiting time for users to obtain their results. Also, the standard deviation of TAT of our method is the smallest, having a reduction of 60.61% over Default. Small standard deviation means TAT of different users will not differ too much, which showcases the fairness of our algorithm. Our QPU time is the second shortest (a bit longer than SNSF), which attains 51.86% reduction over Default. The PST reduction of our method is only 2.47%. Hence, our method can achieve significant improvements in QPU time and TAT at a trivial cost of fidelity.
@@
Compared with other methods, our method is superior in QPU time and TAT. Three greedy baselines share similar TRF and PST with our method. Though QuMC can ensure high PST (positive $\Delta$PST due to randomness), the QPU time and TAT are about twice longer than our method. The improvements of QuMC in time metrics over Default is rather limited.
4.3. Results on a Quantum Computer
@@
As shown in Tbl. 1, the performance of our method is even better on the quantum processor Xiaohong. We still achieve the shortest average TAT, significantly decreasing TAT of Default by 93.2%. SNSF achieves the shortest QPU time as expected, while the maximum and standard deviation of TAT of ESTF are the lowest, indicating the fairness of ESTF. Among all methods, our method has the second lowest QPU time (81.76% reduction over Default) and standard deviation of TAT (87.31% reduction over Default). Hence, our method can significantly reduce time overhead for both users and suppliers, and meanwhile ensure enough fairness.
@@
The superiority of our method on Xiaohong owes to the large TRF (7.18) and our queue rearranging. With 66 physical qubits, Xiaohong can accommodate more programs in each execution. Hence, the execution times are largely diminished compared to Default, resulting in shorter QPU time and TAT. In addition, our priority score can perceive the potential influence of each job on time metrics, arranging those highly influential programs to the head of the queue. Hence, QPU time and TAT are further reduced.
PST on Xiaohong is much lower in general, because noise on Xiaohong is more severe than on the noise model. Our method has 5.17% reduction in PST over Default, slightly bigger than QuMC (3.81%). However, TRF of QuMC is only 1.98. Hence, its QPU Time and average TAT are 3.4$\times$ and 6.0$\times$ longer than ours. Huge growth in execution latency for a small improvement in fidelity is unworthy.
@@ 4.4. Runtime Analysis
As shown in Tbl. 1, the runtime (RT) of our method is close to Default and three greedy baselines. Since routing occupies most of the runtime, our method will not introduce much additional overhead. QuMC costs much more time than us, especially on Xiaohong, because it will repeat routing if a program is unsuitable for its strategy. The runtime of QuMC increases dramatically with the growth in the number of physcal qubits, while our method avoids this issue, indicating the scalability of our method. Note that our scheduling algorithm can run in a pipelining manner with the circuit execution. Hence, it will not affect the running of QPU.
4.5. Sensitivity Analysis

### v2

@@ 3.2.2. Qubit Partitioning
After rearranging the updated queue, we need to select a number of jobs to be executed in parallel in this iteration. Selected jobs must share no common physical qubits with each other, so we should partition the coupling graph into separate parts. Concretely, we pick out the first job in the sorted queue to conduct qubit partitioning. If the partitioning algorithm cannot find a valid partition, it will skip the job, and proceed to the next. Otherwise, we will go on to the qubit mapping step. We use the qubit partitioning algorithm introduced in (Niu and Todri-Sanial, 2023), because it considers both the noise and topology of the QPU, and substantially reduces the search space by limiting the starting points.
This method chooses physical qubits with higher degrees than the largest logical degree as starting points. If such qubits do not exist, it chooses physical qubits with the highest degree as starting points. Then, it adds a neighboring qubit with the highest fidelity degree to the partition iteratively until the number of selected physical qubits equals that of logical qubits. The fidelity degree $D_{f}^{(i)}$ is calculated as $D_{f}^{(i)}=2\times\sum\nolimits_{j\in N(i)}r_{2q}^{(i,j)}+r_{ro}^{(i)}$, where $N(i)$ denotes the neighboring qubits of $Q_{i}$. $r_{2q}^{(i,j)}$ and $r_{ro}^{(i)}$ are the reliability of two-qubit gates on edge $(Q_{i},Q_{j})$ and measurements (readout) on $Q_{i}$. Finally, the partition with the best fidelity score is selected. The fidelity score $S_{f}$ is derived from Eq.(2):
(2) $S_{f}=-N_{2q}\times(1-\overline{r}_{2q})-N_{ro}\times(1-\overline{r}_{ro}),$
where $\overline{r}_{2q}$ and $\overline{r}_{ro}$ are average reliability of two-qubit gates and measurements in this partition. $N_{2q}$ and $N_{ro}$ is the number of the two operations.
@@ 3.2.3. Qubit Mapping
(3) $EPST^{*}=\prod_{i=1}^{N_{1q}}r_{1q}^{(o_{i})}\cdot\prod_{i=1}^{N_{2q}}r_{2q}^{(d_{i})}\cdot\prod_{i=1}^{N_{ro}}r_{ro}^{(m_{i})}\cdot\prod_{i=1}^{n}r_{a}^{(i)}\cdot\prod_{i=1}^{n}r_{p}^{(i)},$
where $r_{1q}^{(*)}$, $r_{2q}^{(*)}$, and $r_{ro}^{(*)}$ denote the reliability of one-qubit gates, two-qubit gates, and measurements. $N_{1q}$, $N_{2q}$, and $N_{ro}$ is the number of the three operations. $o_{i}$, $d_{i}$, and $m_{i}$ map operations to their locations in the circuit. Since our $EPST^{*}$ score considers each gate’s reliability separately, it is more accurate than $EPST$, especially given high variance of reliability. Moreover, $r_{a}^{(i)}$ and $r_{p}^{(i)}$, the probability of amplitude damping error and phase damping error not occurring on the i-th qubit, are involved in $EPST^{*}$ to perceive the impact of decoherence. They can be calculated as: $r_{a}^{(i)}=\exp{(-t_{c}/T_{1}^{(i)})}$, $r_{p}^{(i)}=\exp{(-t_{c}/T_{\phi}^{(i)})}$, $T_{\phi}^{(i)}=T_{1}^{(i)}T_{2}^{(i)}/(2T_{1}^{(i)}-T_{2}^{(i)})$. $T_{1}^{(i)}$ and $T_{2}^{(i)}$ represent the relaxation time and dephasing time of the i-th qubit. The circuit time $t_{c}$ can be calculated by traversing the DAG of a circuit in the topological order as described in Alg. 1. We integrate $EPST^{*}$ in our noise-aware initial mapping algorithm in Alg. 2.
3.3. Complexity Analysis
Given the number of jobs $K$, the number of gates $g$, the number of physical qubits $N$, the number of logical qubits $n$ ($n<N$), the number of starting points $m$ ($m<N$), and the number of repeats $r$, we can calculate the time complexity of our method. The complexity of qubit partitioning and routing is $O\big{(}mn^{2}+N\log(N)+g\big{)}$ and $O(gN^{2.5})$, respectively (Niu and Todri-Sanial, 2023).
Queue rearranging. Calculating the priority score takes $O(K)$ time. The main overhead of queue rearranging lies in sorting the queue, which takes $O\big{(}K\log(K)\big{)}$ time. Hence, the complexity of queue rearranging is $O\big{(}K\log(K)\big{)}$.
Since every job should undergo qubit partitioning, initial mapping, and routing, the total complexity is $O(KrgN^{2.5}+Krn+Kmn^{2})$. Therefore, the overall time complexity is $O\big{(}K\log(K)+KrgN^{2.5}+Krn+Kmn^{2}\big{)}$. In normal circumstances, the number of repeats $r$ is a small constant and we have $g>m$, so the complexity can be reduced to $O\big{(}K\log(K)+KgN^{2.5}\big{)}$. The routing overhead $O(KgN^{2.5})$ is the dominant part, which is unavoidable. Queue rearranging only incurs trivial overhead compared with routing.
35
@@ 46             $Best\_initial\_mapping=Initial\_mapping$
4. Experiments
4.1. Protocols
Dataset. We construct our dataset from RevLib [^2] RevLib can be accessed at https://www.revlib.org/. It contains quantum circuits realizing specific gates like a Toffoli gate, arithmetic functions like a 1-bit adder, etc. (Soeken et al., 2012), a benchmark of reversible and quantum circuits, which is widely used in related works (Li et al., 2019; Liu and Dou, 2021; Niu and Todri-Sanial, 2023). Circuits with extremely large width or depth are unsuitable for the noise model and real quantum hardware, so we filter the data. First, we choose circuits with width no more than 16 to be suitable for the 16-qubit noise model. Second, we translate circuits to fit the basis gate set of the noise model and Xiaohong. Third, we choose translated circuits with depth smaller than 100 to guarantee relatively high fidelity. Finally, the number of candidate circuits on the noise model and Xiaohong is 77 and 20, respectively. The statistics on our dataset is listed in Tbl. 1.
Table 1. Statistics on our dataset.
Environment $N_{2q}$Gates (g) Depth Width
range avg range avg range avg range avg
Noise Model [5, 156] 26.2 [7, 391] 70.8 [5, 99] 44.1 [3, 16] 6.0
Xiaohong [5, 22] 13.6 [19, 111] 66.6 [15, 68] 44.1 [3, 16] 6.7
Then, we sample from candidate circuits to construct our dataset. We focus on the congestion scene, where there are some initial jobs and much more jobs to be submitted. Due to the limitation of quantum resources and time, the number of initial jobs is 44 on average, and the number of new submitted jobs is 400. The submission time $t_{i}$ of initial jobs is set as 0. For new submitted jobs, $t_{i+1}\in\{t_{i},t_{i}+1\}$. According to our observation, at peak periods on IBM quantum cloud, there are approximately two jobs submitted per second on average. Hence, the ratio is in line with the congestion in reality. For the noise model, the number of shots $s_{i}$ in each job ($c_{i},s_{i},t_{i}$) is set as a random integer from 1K to 20K. For Xiaohong (QuantumCTek), we modify the range of $s_{i}$ as $\left[500,10K\right]$ to reduce running overhead. The length of dataset is 10.
Table 2. Performance comparison between different methods (with best in bold and second best underlined.
@@ 4.1. Protocols
Noise Model (Guadalupe) FIFO 925.94 0 5155 2591 0 1484 0 197 1 72.88
FIFO-p 502.90 -45.69 2323 1070 -58.71 610 -58.88 208 2.43 70.06
NAQJS 443.76 -50.07 2280 802 -69.03 585 -60.61 202 2.41 69.65
QuMC 734.68 -20.66 3819 2021 -21.99 1046 -29.52 296 1.41 72.87
QuCloud 601.62 -35.03 2897 1335 -48.47 785 -47.11 733 1.89 71.10
Noise Model (Chain) FIFO 925.94 0 5155 2591 0 1484 0 - 1 69.55
FIFO-p 529.32 -42.83 2273 1073 -58.57 632 -57.39 - 2.30 67.68
NAQJS 516.40 -44.23 2413 876 -66.21 663 -55.35 - 2.23 67.98
QuMC 683.17 -26.22 3405 1611 -37.84 936 -36.90 - 1.60 69.66
QuCloud 721.09 -22.12 3688 1873 -27.70 1052 -29.14 - 1.46 69.90
Xiaohong FIFO 468.31 0 4733 2377 0 1361 0 110 1 45.86
FIFO-p 95.89 -79.52 455 217 -90.88 118 -91.33 112 8.27 32.31
NAQJS 96.75 -79.34 688 156 -93.45 176 -87.09 133 6.62 35.70
NAQJS ($\eta=2/7$) 216.21 -53.83 1826 594 -74.99 484 -64.41 125 2.53 42.48
QuMC 270.78 -42.18 2177 917 -61.43 627 -53.90 515 2.31 43.33
QuCloud 372.49 -20.46 3637 2025 -14.82 1191 -12.45 1526 1.34 43.97
@@
  • $\dagger$: our proposed method. $\Delta$ QPU Time(%): percentage difference to the QPU time of FIFO. TAT[s]: turnaround time in seconds. RT[s]: runtime of scheduling algorithms in seconds. $\Delta$ avg(%): percentage difference to the average of FIFO. $\Delta$ std(%): percentage difference to the standard deviation of FIFO. TRF: Trial Reduction Factor (Das et al., 2019).
Baselines. As there are few existing methods addressing the scheduling problem proposed in this paper, we devise four baselines (i.e. FIFO, FIFO-p, QuMC, QuCloud) to verify the effectiveness of NAQJS. First-In-First-Out (FIFO) denotes the current running mode of quantum computers. Specifically, each submitted quantum job is executed in serial according to their submission time. FIFO-p represents that all quantum jobs are executed in parallel according to their submission time. In each execution round, quantum jobs will be allocated on the quantum processor in chronological order until the next job cannot be accommodated. The partition and mapping methods of FIFO and FIFO-p are the same as NAQJS. Additionally, we adapt QuMC (Niu and Todri-Sanial, 2023) and QuCloud (Liu and Dou, 2021) to fit QJSP by merging their queuing method into our framework. For baselines, we do not explicitly restrain the maximum usage of physical qubits.
Parameter Setting and Experiment Environment. Experiments are performed on the IBM Guadalupe noise model and its chain version (discussed in Sec. 4.5) as simulation, and the physical Xiaohong (QuantumCTek) quantum processor. For Xiaohong, the average relaxation time $\overline{T}_{1}$ and $\overline{T}_{2}$ are 27.35 $\mu s$ and 20 $\mu s$. The average reliability of one-qubit gate $\overline{r}_{1q}$, two-qubit gate $\overline{r}_{2q}$, and measurement $\overline{r}_{ro}$ are 99.85%, 97.07%, and 93.97%.
@@
4.2. Results on the Noise Model Guadalupe
@@
As shown in Tbl. 2, our method NAQJS achieves the shortest average turnaround time (TAT) across all methods. It reduces TAT of FIFO by nearly 70 percent, which will significantly cut down the waiting time for users to obtain their results. Also, the standard deviation of TAT of NAQJS is the smallest, having a reduction of 60.61% over FIFO. Small standard deviation means TAT of different users will not differ too much, which showcases the fairness of NAQJS. The maximum TAT of NAQJS ranks second (only a bit longer than FIFO-p), indicating that no job will wait too long to be executed, further strengthening the fairness of our method. Besides, our QPU time is the shortest among all the five methods, which attains 50.07% reduction over FIFO. The PST reduction of NAQJS is only 3.23%. In other words, NAQJS can achieve significant improvements in QPU time and TAT at a trivial cost of fidelity.
@@
Compared with other methods, NAQJS is superior in QPU time and TAT. Though QuMC can ensure high PST, the QPU time and TAT are about twice longer than NAQJS. The improvements of QuMC in time metrics over FIFO is rather limited.
4.3. Results on Xiaohong Quantum Processor
@@
As shown in Tbl. 2, we still achieve the shortest average TAT, significantly decreasing TAT of FIFO by 93.45%. Among all methods, NAQJS has the second lowest QPU time (79.34% reduction over FIFO) and standard deviation of TAT (87.09% reduction over FIFO). Hence, NAQJS can significantly reduce time overhead for both users and suppliers, and meanwhile ensure enough fairness.
@@
The superiority of NAQJS on Xiaohong owes to the large TRF (6.62) (Das et al., 2019) and our queue rearranging. TRF is the ratio of the number of trials when circuits run in series to that when circuits run in parallel. With 66 physical qubits, Xiaohong can accommodate more jobs in each execution. Hence, the execution times are largely diminished compared to FIFO, resulting in shorter QPU time and TAT. In addition, our priority score can perceive the potential influence of each job on time metrics, arranging those highly influential jobs to the head of the queue. Hence, QPU time and TAT are further reduced. Though FIFO-p slightly outperforms NAQJS in QPU time by 0.18 percent due to larger TRF (8.27), its average TAT is 2.57 percent larger than ours, and its PST is 3.39 percent lower than us. Considering all these metrics, NAQJS performs the best in general.
PST on Xiaohong is much lower, because noise on Xiaohong is more severe than on the noise model. NAQJS has 10.16% reduction in PST over FIFO while QuMC and QuCloud keep relative high PST (43.33% and 43.97%). However, their TRF is only 2.31 and 1.34. Hence, their QPU Time and average TAT are more than twice longer than ours. We further validate that NAQJS surpasses QuMC and QuCloud in time metrics by a large gap even when PST is close. When we set the maximum usage $\eta$ as 2/7 in Tbl. 2, the PST (42.48%) is almost the same as QuMC and QuCloud, but the QPU time and average TAT are still much lower than theirs (over 10%).
@@ 4.4. Runtime Analysis
As shown in Tbl. 2, the runtime (RT) of NAQJS is close to FIFO. As routing occupies most of the runtime, NAQJS will not introduce much additional overhead. QuMC and QuCloud cost much more time than us, especially on Xiaohong. QuMC and QuCloud will repeat routing if a job is unsuitable for their strategy. The runtime of them increases dramatically with the growth in the number of physcal qubits, while NAQJS avoids this issue, indicating the scalability. Note that NAQJS can run in a pipelining manner with the circuit execution, and the QPU time plus the total extra time (2279.76 s for Guadalupe and 780.09 s for Xiaohong) is far longer than our runtime. Hence, it will not affect the running of QPU.
4.5. Impact of QPU Topology
Quantum processors may have different topology, leading to distinct connectivity of qubits. The topology of the noise model Guadalupe is ring-type shown in Fig. 1b. To explore the impact of QPU Topology on QJSP metrics, we disconnect $Q_{1}$ from $Q_{4}$ in Guadalupe and obtain chain-type topology depicted in Fig. 5. The noise information of every qubit remains unchanged. The results are listed in Tbl. 2. The connectivity of chain-type topology (Chain) is worse than ring-type topology (Guadalupe), so the performance of all the methods deteriorates. NAQJS still outperforms others in both QPU time and average TAT. Besides, the PST gap between our method and FIFO gets smaller.
Figure 5. Chain-type topology.
Figure 6. (a) Impact of $\eta$ on average TAT and PST on Xiaohong. (b) Impact of $\gamma$ on average TAT and standard deviation of TAT on the noise model. (c) Impact of $\Delta t$ on average TAT and maximum TAT on the noise model.
Figure 7. (a) Impact of $\alpha$ and $\beta$ on average TAT on the noise model. (b) Impact of $\alpha$ and $\beta$ on QPU time on the noise model. (c) Impact of noise level on PST (solid lines) and average TAT (dashed lines) on the noise model.
4.6. Impact of Noise Level
The noise level can affect the metrics in QJSP, especially PST. To investigate this, we multiply the noise of one-qubit gates, two-qubit gates and measurements by the noise level (0.5, 1, or 2). As shown in Fig. 7c, PST of all the methods drops drastically with the increase of the noise level. QuMC and QuCloud sacrifice average TAT for PST when noise level is large. QuMC almost degrades to the serial running mode (one job per execution) when noise level reaches 2. By contrast, average TAT of NAQJS is insensitive to the noise level. Therefore, NAQJS can keep low average TAT and adequately high PST even when the noise condition is poor.
More importantly, the PST gap between NAQJS and other methods narrows when noise level decreases. Researchers are currently devoted to manufacturing larger-scale quantum processors and fabricating noiseless logical qubits. As a result, users will pay more and more attention to time metrics in the future. The advantage in time metrics of NAQJS will be further amplified with the increase in qubit number and decrease in noise.
4.7. Sensitivity Analysis of Hyperparameters